Assume that Josie forms JES as a C-Corporation.

•JES earns taxable income of $300,000.

•JES will distributes 40% of its after-tax earnings annually as a dividend.

•Josie’s marginal ordinary tax rate is 37% and her dividend tax rate is 23.8% (including the net investment income tax).

What is the overall tax rate on JES’s taxable income?
Answer #1

Overall tax rate:

= (total income * marginal tax rate + income left after marginal tax*dividend percentage * tax on dividend rate)*100

   = (1 * 0.37 + 0.63 *0.4*0.238)*100

    = (0.37 + 0.0598)*100

   = 0.4298 * 100

= 42.98 % Answer

Please upvote if this helps.